OpenStudy (anonymous):

4. How many molecules of H2O are there in 1.0 g of H2O? Please help me how to do this? :O

OpenStudy (anonymous):

The molar weight of H2O is 18, because Hydrogen weighs 1g/mol, and Oxygen weighs 16g/mol. Therefore, to find the number of moles of H2O in one gram, divide the given amount by the molar weight. \[\frac{ 1 }{ 18 }mol\]

OpenStudy (anonymous):

Grams to moles to molecules. 1.0 g H20 * 1 mol/18.016 grams H20 * 6.022 X 10^23 molecules/1 mol H20 = number of molecules in 1 gram of H20 Does that make sense? :)

thomaster (thomaster):

First calculate the amount of mol H2O. H=1.00794 g/mol O=15.9994 g/mol So the molecular weight of H2O = (2*1.00794)+15.9994= 18.01528 g/mol 18.01528 grams = 1 mol. So 1 gram = 1/18.01528 = 0.0555 mol H2O Multiply that by Avogadro's number ( 6.0221413*10^23 ) 0.0555 * 6.0221413*10^23 = the number of molecules in 1 gram H2O